OpenStudy (anonymous):

exact value- answer must be either an integer or franction. sin(2/3 pi)= cos(2/3pi)=

OpenStudy (anonymous):

sin(2/3 pi)= sqrt(3)/2 cos(2/3pi)= -1/2

OpenStudy (anonymous):

I think you should understand the quadrant's rule :)